OpenStudy (anonymous):

Please help! Step by step please, and fully simplified. Simplify the expression. cos^2 x + sin^2 x / cot^2 x - csc^2 x

OpenStudy (campbell_st):

well you should know sin^2 + cos^2 = 1 and if you divide every term by sin^2 you get 1 + cot^2 = csc^2 rewriting it to the form you need 1 = csc^2 - cot^2 make the necessary substitutions to get the simplified answer. hope it helps

OpenStudy (anonymous):

Thank you! But what kind of substitutions would I make to get the simplified answer?

OpenStudy (campbell_st):

well \[\frac{\sin^2 + \cos^2}{\csc^2 - \cot^2} = \frac{1}{-1(\cot^2 - \csc^2)} = \frac{1}{-1 \times 1} = -1\]